OpenStudy (anonymous):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B 1/2 is .

OpenStudy (anonymous):

simplify this

OpenStudy (anonymous):

number cubes!!!

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

think you are missing something

OpenStudy (anonymous):

u r missing somewhat some thing

OpenStudy (anonymous):



OpenStudy (anonymous):

i remember when these where called dice. ask your teacher when they starting being called "number cubes"!!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

remember when these where called dice. ask your teacher when they starting being called number cubes

OpenStudy (anonymous):

what is the actual question, the cut and paste failed

OpenStudy (anonymous):

simplify the fraction

OpenStudy (anonymous):

then part b says . What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?

OpenStudy (anonymous):

are two dice being rolled?

OpenStudy (king):

part b.1/4 or 1.....i thinks....

OpenStudy (anonymous):

one land 2,3,4, other lands 1,5,6 probability is \[\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\]

OpenStudy (anonymous):

whats part to simplify that fraction

OpenStudy (anonymous):

i can't read it. use the equation editor

OpenStudy (anonymous):

\[(1/2 a+1/2 b)^{2}\]

OpenStudy (anonymous):

@satellite73 there it is

OpenStudy (anonymous):

is it \[(\frac{1}{2}a+\frac{1}{2}b)^2\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so you get \[\frac{1}{4}a^2+\frac{1}{2}ab+\frac{1}{4}b^2\]

OpenStudy (anonymous):

u simplified

OpenStudy (anonymous):

@satellite73